# Joint Pedestrian and Vehicle Traffic Optimization in Urban Environments using Reinforcement Learning

### v1

@@ IV-B2 Signalized
Figure 4: (a) Vehicle-pedestrian right-of-way conflicts in unsignalized mid-block crosswalks increase substantially with traffic demand, from $1.8$ conflicts at $0.5$x demand to $28.1$ at $2.5$x demand. The error bars show standard deviation values rising at higher demand levels, indicating safety outcomes become both worse and more unpredictable as traffic volumes increase. (b) Emergence of traffic signal coordination during training. Average number of mid-block traffic lights simultaneously set to green for vehicular flow shows an upward trend (from approximately $3.55$ to $3.75$) indicating that the RL agent gradually learns to coordinate multiple signals. (c) Total signal phase switches after the warmup period across all signals in the network. RL agent demonstrates adaptive switching, with more than five fold increase in switching frequency compared to fixed-time signalized control. Plots (a) and (c) data are averaged over $10$ runs.
@@ IV-C Results
Pedestrian Wait Times. At $1$x demand, our approach achieves an average wait time per pedestrian of $2.1$ seconds compared to $5.9$ seconds for Signalized and $2.1$ seconds for Unsignalized. This represents a $65\%$ reduction compared to Signalized while matching Unsignalized. As demand increases, our approach maintains its efficiency: at $2$x demand, $2.0$ seconds versus $6.0$ seconds for Signalized ($67\%$ reduction) and $2.2$ seconds for Unsignalized. The total pedestrian wait time shows a similar pattern (upto $67\%$ reduction), with our approach consistently outperforming both Signalized and Unsignalized across the demands.
Vehicle Wait Times. At $1$x demand, our approach achieves an average wait time per vehicle of $20.8$ seconds compared to $31.3$ seconds for Signalized and $20.7$ seconds for Unsignalized. This represents a $34\%$ reduction compared to Signalized while maintaining comparable performance to Unsignalized. This pattern continues at higher demands: at $2$x demand, $20.5$ seconds versus $42.6$ seconds for Signalized ($52\%$ reduction) and $25.9$ seconds for Unsignalized. The total accumulated vehicle wait time at $2$x demand shows similar improvements: $1.16$ hours for Signalized, $0.55$ hours for our approach, and $0.56$ hours for Unsignalized.
@@
Signal Coordination. Our policy autonomously learns to coordinate mid-block crosswalks, similar to the “green wave” effect observed in other RL traffic systems [51, 52]. The action space models each mid-block signal independently using Bernoulli distribution (as described in Section III), with no built-in mechanism for coordination. Yet as shown in Figure 4(b), the RL agent learns to increase the number of mid-block traffic signals simultaneously set to vehicle green phase. The trend line shows this coordination improve from approximately $3.55$ to $3.75$ synchronized green signals over the course of training—a significant shift when considering each data point represents an average over $1440$ actions taken across $24$ parallel actors. The coordination of signal timing reduces the number of stops and enables more efficient vehicular flow through the corridor.
@@ V Conclusion and Future Work
Under high demands, qualitative analysis of the rollouts reveal the propagation of vehicle queues upstream, when signals are located close to each other (back-spill effect). This may have resulted from our modeling assumption that each signal operates independently. There exist many future directions. First, the framework could explicitly incorporate correlations between adjacent signals to address this limitation. Second, with the rise of self-driving technology, we would like to incorporate mixed traffic control [53, 54, 55, 56, 57, 58] into our framework. Finally, the framework could explore other topics such as adversarial robustness evaluation [59] and effect of real-world traffic disruptions [60, 61].
@@ Acknowledgments
This research is supported by NSF IIS-$2153426$. The authors thank NVIDIA and the Tickle College of Engineering at University of Tennessee, Knoxville for their support.

### v2

@@ IV-B2 Signalized
Figure 4: (a) Vehicle-pedestrian right-of-way conflicts in unsignalized mid-block crosswalks increase substantially with traffic demand, from $1.8$ conflicts at $0.5$x demand to $28.1$ at $2.5$x demand. The error bars show standard deviation values rising at higher demand levels, indicating safety outcomes become both worse and more unpredictable as traffic volumes increase. (b) Emergence of traffic signal coordination behavior during training. The average number of mid-block traffic lights simultaneously set to green for vehicular flow shows an upward trend (from approximately $3.55$ to $3.75$) indicating that the RL agent gradually learns to coordinate multiple signals. (c) Total signal phase switches after the warmup period across all signals in the network. RL agent demonstrates adaptive switching, with more than fivefold increase in switching frequency compared to fixed-time signalized control. Plots (a) and (c) data are averaged over $10$ runs.
@@ IV-C Results
Pedestrian Wait Times. At $1$x demand, our approach achieves an average wait time per pedestrian of $2.1$ seconds compared to $5.9$ seconds for Signalized and $2.1$ seconds for Unsignalized. This represents a $65\%$ reduction compared to Signalized while matching Unsignalized. As demand increases, our approach maintains its efficiency: at $2$x demand, $2.0$ seconds versus $6.0$ seconds for Signalized ($67\%$ reduction) and $2.2$ seconds for Unsignalized. The total pedestrian wait time shows a similar pattern (up to $67\%$ reduction), with our approach consistently outperforming both Signalized and Unsignalized across the demands. Even as pedestrian demand increases, our agent consistently maintains efficient crossing opportunities without compromising safety (no pedestrian-vehicle conflicts).
Vehicle Wait Times. At $1$x demand, our approach achieves an average wait time per vehicle of $20.8$ seconds compared to $31.3$ seconds for Signalized and $20.7$ seconds for Unsignalized. This represents a $34\%$ reduction compared to Signalized while maintaining comparable performance to Unsignalized. This pattern continues at higher demands: at $2$x demand, $20.5$ seconds versus $42.6$ seconds for Signalized ($52\%$ reduction) and $25.9$ seconds for Unsignalized. The total vehicle wait time at $2$x demand shows similar improvements: $1.16$ hours for Signalized, $0.55$ hours for our approach, and $0.56$ hours for Unsignalized.
@@
Signal Coordination. Our policy autonomously learns to coordinate mid-block crosswalks, similar to the “green wave” effect observed in other RL traffic systems [53, 54]. The action space models each mid-block signal independently using a Bernoulli distribution (as described in Section III), with no built-in mechanism for coordination. Yet as shown in Figure 4(b), the RL agent learns to increase the number of mid-block traffic signals simultaneously set to vehicle green phase. The trend line shows this coordination increasing from approximately $3.55$ to $3.75$ synchronized green signals over the course of training—a significant shift when considering each data point represents an average over $1440$ actions taken across $24$ parallel actors. The coordination of signal timing reduces the number of stops and enables more efficient vehicular flow through the corridor.
@@ V Conclusion and Future Work
While our approach demonstrates significant improvements, several limitations exist. First, under high demand, we observe vehicle queue propagation upstream when signals are closely spaced (a back-spill effect). This is likely a result of our modeling assumption, i.e., we assume signals operate independently. Second, the high switching frequency exhibited by our policy, while reducing wait times, breaks traffic flow continuity and may increase vehicle energy consumption because of more frequent stop-and-go events. Third, our evaluation was limited to a university campus setting with its specific pedestrian and vehicle patterns; generalization to heterogeneous traffic conditions, different corridor geometries including irregular intersections, and diverse urban environments remains to be validated. Fourth, while our baselines include common real-world approaches (fixed-time control and unsignalized crosswalks), direct comparison with other RL-based controllers proved challenging: existing RL methods either exclude pedestrian considerations entirely or require separate agents at each intersection with distributed training infrastructure and communication protocols, whereas our approach uses a single policy controlling all eight signals through centralized training.
Future work could address these limitations by explicitly modeling correlations between adjacent signals, incorporating traffic flow continuity or energy consumption metrics into the reward function, and evaluating performance across diverse urban settings with varying pedestrian-vehicle ratios and geometric configurations. Additional future directions include mixed-traffic control [55, 56], adversarial robustness evaluation [57], and real-world traffic disruption analysis [58, 59].
@@ Acknowledgments
This research is supported by NSF IIS-$2153426$ and ECCS-$2332210$. The authors also thank NVIDIA and the Tickle College of Engineering at University of Tennessee, Knoxville for their support.